### v1

# Recursive variational Gaussian approximation with the Whittle likelihood for linear non-Gaussian state space models

Abstract
Parameter inference for linear and non-Gaussian state space models is challenging because the likelihood function contains an intractable integral over the latent state variables. Exact inference using Markov chain Monte Carlo is computationally expensive, particularly for long time series data. Variational Bayes methods are useful when exact inference is infeasible. These methods approximate the posterior density of the parameters by a simple and tractable distribution found through optimisation. In this paper, we propose a novel sequential variational Bayes approach that makes use of the Whittle likelihood for computationally efficient parameter inference in this class of state space models. Our algorithm, which we call Recursive Variational Gaussian Approximation with the Whittle Likelihood (R-VGA-Whittle), updates the variational parameters by processing data in the frequency domain. At each iteration, R-VGA-Whittle requires the gradient and Hessian of the Whittle log-likelihood, which are available in closed form for a wide class of models. Through several examples using a linear Gaussian state space model and a univariate/bivariate non-Gaussian stochastic volatility model, we show that R-VGA-Whittle provides good approximations to posterior distributions of the parameters and is very computationally efficient when compared to asymptotically exact methods such as Hamiltonian Monte Carlo.
Keywords: Autoregressive process; Fast Fourier transform; Hamiltonian Monte Carlo; Stationary time series; Stochastic volatility models.
1 Introduction
State space models (SSMs) are a widely used class of hierarchical time series models employed in a diverse range of applications in fields such as epidemiology (e.g., Dukic et al.,, 2012), economics (e.g., Chan and Strachan,, 2023), finance (e.g., Kim et al.,, 1998), and engineering (e.g., Gordon et al.,, 1993). In an SSM, the observed data depend on unobserved, time-evolving latent state variables and unknown parameters. The observations are assumed to be independent given the latent states, the dynamics of the latent states are described through a Markov process, and often a simplifying assumption of linearity is made. Bayesian inference for parameters of linear non-Gaussian state space models is challenging since the likelihood function involves an intractable integral over the latent states.
A popular class of inference methods for non-Gaussian SSMs is Markov chain Monte Carlo (MCMC). Popular MCMC methods include particle MCMC (Andrieu et al.,, 2010) and Hamiltonian Monte Carlo (HMC, e.g., Neal,, 2011; Betancourt and Girolami,, 2015). While MCMC methods can provide samples from the exact posterior distribution, they tend to be computationally expensive.
Variational Bayes (VB) methods are becoming increasingly popular for Bayesian inference for state space models (Zammit-Mangion et al.,, 2012; Gunawan et al.,, 2021; Frazier et al.,, 2023). VB approximates the posterior distribution of model parameters with a simpler and more tractable class of distributions constructed using so-called variational parameters, which are unknown and estimated via optimisation techniques. While VB methods tend to be more computationally efficient than MCMC methods, they provide approximate posterior distributions rather than samples from the exact posterior distributions. For a review of VB methods, see Blei et al., (2017).
There are two main classes of VB methods: batch and sequential. A batch VB framework (Ong et al.,, 2018; Tan and Nott,, 2018; Gunawan et al.,, 2024) requires an optimisation algorithm to repeatedly process the entire dataset until variational parameters converge to a fixed point. A sequential VB framework, on the other hand, updates variational parameters using one observation at a time. Sequential methods only pass through the entire dataset once, and are thus ideal for settings with large datasets. Sequential VB methods are available for statistical models for independent data (Lambert et al.,, 2022), for classical autoregressive models for time-series data (Tomasetti et al.,, 2022), for generalised linear mixed models (Vu et al.,, 2024), and for SSMs where the state evolution model is linear and Gaussian (Beal,, 2003, Chapter 5; Zammit Mangion et al.,, 2011; Mangion et al.,, 2011). Campbell et al., (2021) proposed a sequential VB approach that is applicable to an SSM with non-linear state dynamics and non-Gaussian measurement errors, but this approach can be computationally expensive as it requires solving an optimisation problem and a regression task for each time step.
Our article proposes a sequential VB approach to estimate parameters in linear non-Gaussian state space models. This approach builds on the Recursive Variational Gaussian Approximation for Latent variable models (R-VGAL) algorithm of Vu et al., (2024). We first transform time-domain data into frequency-domain data and then use the Whittle likelihood (Whittle,, 1953) in a sequential VB scheme. The logarithm of the Whittle likelihood function is a sum of the individual log-likelihood components in Fourier space, and is thus well-suited for use in sequential updating. We call our algorithm the Recursive Variational Gaussian Approximation with the Whittle likelihood (R-VGA-Whittle) algorithm. At each iteration, R-VGA-Whittle sequentially updates the variational parameters by processing one frequency or a block of frequencies at a time. Although sequential in nature, and requiring only one pass through the frequencies, we note that the R-VGA-Whittle algorithm requires the full dataset to be available in order to construct the periodogram, from which we obtain the frequency-domain data.
We demonstrate the performance of R-VGA-Whittle in simulation studies involving a linear Gaussian state space model and univariate and bivariate non-Gaussian stochastic volatility models. We show that R-VGA-Whittle is accurate and is much faster than Hamiltonian Monte Carlo with the exact likelihood (HMC, Duane et al.,, 1987; Neal,, 2011; Hoffman and Gelman,, 2014) as well as HMC based on the Whittle likelihood; hereafter, we refer to these two methods as HMC-exact and HMC-Whittle, respectively. We subsequently apply R-VGA-Whittle to analyse exchange rate data.
The rest of the article is organised as follows. Section 2 gives an overview of SSMs and the univariate and multivariate Whittle likelihood functions, and then introduces the R-VGA-Whittle and HMC-Whittle algorithms. Section 3 compares the performance of R-VGA-Whittle, HMC-Whittle and HMC with the exact likelihood in several examples involving linear Gaussian and non-Gaussian SSMs. Section 4 summarises and discusses our main findings. The article also has an online supplement with additional technical details and empirical results.
2 Methodology
We give an overview of SSMs in Section 2.1, and the univariate and multivariate Whittle likelihoods in Section 2.2. Section 2.3 introduces R-VGA-Whittle, while Section 2.4 proposes a blocking approach to improve computational efficiency. Section 2.5 gives details of HMC-Whittle.
2.1 State space models
An SSM consists of two sub-models: a data model that describes the relationship between the observations and the latent state variables, and a state evolution model that describes the (temporal) evolution of the latent states.
In SSMs, the latent state variables $\{\mathbf{x}_{t}\}_{t=1}^{T}$ are characterised by an initial density ${\mathbf{X}_{1}\sim p(\mathbf{x}_{1}\mid{\bm{\theta}})}$ and state transition densities
$\mathbf{X}_{t}\mid(\mathbf{X}_{t-1}=\mathbf{x}_{t-1})\sim p(\mathbf{x}_{t}\mid\mathbf{x}_{t-1},\bm{\theta}),\quad t=2,\dots,T,$ (1)
where $\bm{\theta}$ is a vector of model parameters. The observations $\{\mathbf{y}_{t}\}_{t=1}^{T}$ are related to the latent states via
$\mathbf{Y}_{t}\mid(\mathbf{X}_{t}=\mathbf{x}_{t})\sim p(\mathbf{y}_{t}\mid\mathbf{x}_{t},\bm{\theta}),\quad t=1,\dots,T,$ (2)
where each $\mathbf{Y}_{t}$ is assumed to be conditionally independent of the other observations, when conditioned on the latent state $\mathbf{x}_{t}$. A linear SSM is one where $\mathbf{X}_{t}$ and $\mathbf{Y}_{t}$ in (1) and (2) can each be expressed as the sum of two components: a linear function of their conditioning variables and an exogenous random component.
Let $\mathbf{y}_{1:T}\equiv(\mathbf{y}_{1}^{\top},\dots,\mathbf{y}_{T}^{\top})^{\top}$ and $\mathbf{x}_{1:T}\equiv(\mathbf{x}_{1}^{\top},\dots,\mathbf{x}_{T}^{\top})^{\top}$ be vectors of observations and latent states, respectively. The posterior density of the model parameters is $p(\bm{\theta}\mid\mathbf{y}_{1:T})\propto p(\mathbf{y}_{1:T}\mid\bm{\theta})p(\bm{\theta})$, where the likelihood function is
$p(\mathbf{y}_{1:T}\mid\bm{\theta})=\int p(\mathbf{y}_{1}\mid\mathbf{x}_{1},\bm{\theta})p(\mathbf{x}_{1}\mid\bm{\theta})\prod_{t=2}^{T}p(\mathbf{y}_{t}\mid\mathbf{x}_{t},\bm{\theta})p(\mathbf{x}_{t}\mid\mathbf{x}_{t-1},\bm{\theta})\,\mathrm{d}\mathbf{x}_{1:T}.$
This likelihood function is intractable, except for special cases such as the linear Gaussian SSM (e.g., Gibson and Ninness,, 2005). In non-Gaussian SSMs, the likelihood is typically estimated using a particle filter (Gordon et al.,, 1993), such as in the pseudo-marginal Metropolis-Hastings (PMMH) method of Andrieu and Roberts, (2009), which can be computationally expensive for large $T$.
The Whittle likelihood (Whittle,, 1953) is a frequency-domain approximation of the exact likelihood, and offers a computationally efficient alternative for parameter inference in SSMs. Transforming the data to the frequency domain “bypasses” the need for integrating out the states when computing the likelihood, as the Whittle likelihood only depends on the model parameters and not the latent variables. In the following sections, we review the Whittle likelihood and its properties, and then propose a sequential VB approach that uses the Whittle likelihood for parameter inference.
2.2 The Whittle likelihood
The Whittle likelihood has been widely used to estimate the parameters of complex time series or spatial models. It is more computationally efficient to evaluate than the exact likelihood, although it is approximate (e.g., Sykulski et al.,, 2019). Some applications of the Whittle likelihood include parameter estimation with long memory processes (e.g., Robinson,, 1995; Hurvich and Chen,, 2000; Giraitis and Robinson,, 2001; Abadir et al.,, 2007; Shao and Wu, 2007b, ; Giraitis et al.,, 2012), and with spatial models with irregularly spaced data (e.g., Fuentes,, 2007; Matsuda and Yajima,, 2009). The Whittle likelihood is also used for estimating parameters of vector autoregressive moving average (VARMA) models with heavy-tailed errors (She et al.,, 2022), and of vector autoregressive tempered fractionally integrated moving average (VARTFIMA) models (Villani et al.,, 2022).
2.2.1 Univariate Whittle likelihood
Consider a real-valued, zero-mean, second-order stationary, univariate time series $\{y_{t}\}_{t=1}^{T}$ with autocovariance function $C(h;\bm{\theta})\equiv\mathbb{E}(y_{t}y_{t-h})$, where $0\leq h<t$ and $\mathbb{E}(\cdot)$ is the expectation operator. The power spectral density of the time series is the Fourier transform of the autocovariance function (Brillinger,, 2001):
$f(\omega;\bm{\theta})=\sum_{h=-\infty}^{\infty}C(h;\bm{\theta})\exp(-i\omega h),$ (3)
where $\omega\in(-\pi,\pi]$ is the angular frequency and $\bm{\theta}$ is a vector of model parameters. The periodogram, an asymptotically unbiased estimate of the spectral density, is given by
$I(\omega_{k})=\frac{1}{T}\left\lvert J(\omega_{k})\right\rvert^{2},\quad k=-\left\lceil\frac{T}{2}\right\rceil+1,\dots,\left\lfloor\frac{T}{2}\right\rfloor,$ (4)
where $\omega_{k}=\frac{2\pi k}{T}$, and $J(\omega_{k})$ is the discrete Fourier transform (DFT) of $\{y_{t}\}_{t=1}^{T}$:
$J(\omega_{k})=\sum_{t=1}^{T}y_{t}\exp(-i\omega_{k}t).$ (5)
For an observed time series, the Whittle likelihood, denoted by $l_{W}(\bm{\theta})$, is (e.g., Salomone et al.,, 2020)
$l_{W}(\bm{\theta})=-\sum_{k=1}^{\left\lfloor\frac{T-1}{2}\right\rfloor}\left(\log f(\omega_{k};\bm{\theta})+\frac{I(\omega_{k})}{f(\omega_{k};\bm{\theta})}\right).$ (6)
Here, the Whittle likelihood is evaluated by summing over the non-negative frequencies only, since $f(\omega_{k};\bm{\theta})$ and $I(\omega_{k})$ are both symmetric about the origin for real-valued time series. The term $\omega_{k}=0$ is not included since $I(0)=0$ when the time series is demeaned. The term $\omega_{k}=\pi$ is also omitted since it has a different distribution from the other frequencies and its influence is asymptotically negligible (Salomone et al.,, 2020; Villani et al.,, 2022).
Asymptotically, it can be shown that the parameter estimates obtained by maximising the exact likelihood and those obtained by maximising the Whittle likelihood are equivalent (Dzhaparidze and Yaglom,, 1983). The calculation of the Whittle likelihood involves evaluation of the periodogram, which does not depend on the parameters $\bm{\theta}$ and can, therefore, be pre-computed at a cost of $O(T\log T)$ via the Fast Fourier Transform (FFT, Cooley and Tukey,, 1965). Since the periodogram ordinates are asymptotically independent (Whittle,, 1953), the logarithm of the Whittle likelihood is a sum of the individual log-likelihood terms computed over different frequencies.
2.2.2 Multivariate Whittle likelihood
Consider a $d$-variate, real valued, second-order stationary, zero mean time series $\{\mathbf{y}_{t}\}_{t=1}^{T}$ with autocovariance function $\mathbf{C}(h;\bm{\theta})\equiv\mathbb{E}(\mathbf{y}_{t}\mathbf{y}_{t-h}^{\top})$, $0\leq h<t$. At a given angular frequency $\omega$, the power spectral density matrix of this series is
$\mathbf{f}(\omega;\bm{\theta})=\sum_{h=-\infty}^{\infty}\mathbf{C}(h;\bm{\theta})\exp(-i\omega h),\quad\omega\in(-\pi,\pi].$ (7)
The matrix-valued periodogram is
$\mathbf{I}(\omega_{k})=\frac{1}{T}\mathbf{J}(\omega_{k})\mathbf{J}(\omega_{k})^{H},\quad k=-\left\lceil\frac{T}{2}\right\rceil+1,\dots,\left\lfloor\frac{T}{2}\right\rfloor,$ (8)
where $\omega_{k}=\frac{2\pi k}{T}$, $\mathbf{A}^{H}$ denotes the conjugate transpose of a complex matrix $\mathbf{A}$, and $\mathbf{J}(\omega_{k})$, the DFT of the series, is given by
$\mathbf{J}(\omega_{k})=\sum_{t=1}^{T}\mathbf{y}_{t}\exp(-i\omega_{k}t).$ (9)
The Whittle likelihood for the multivariate series $\{\mathbf{y}_{t}\}_{t=1}^{T}$ is (Villani et al.,, 2022):
$l_{W}(\bm{\theta})=-\sum_{k=1}^{\lfloor\frac{T-1}{2}\rfloor}\left(\log\lvert\mathbf{f}(\omega_{k};\bm{\theta})\rvert+\operatorname{Tr}(\mathbf{f}(\omega_{k};\bm{\theta}))^{-1}\mathbf{I}(\omega_{k})\right).$ (10)
2.3 R-VGA-Whittle
Vu et al., (2024) recently extended the Recursive Variational Gaussian Approximation (R-VGA) algorithm of Lambert et al., (2022) to estimate parameters in generalised linear mixed models (GLMMs). Their algorithm, called Recursive Variational Gaussian Approximation for Latent variable models (R-VGAL), is given in Algorithm 1, and described in detail in Section S1 of the online supplement. R-VGAL was successfully applied to grouped data where observations in the same group are correlated. The sequential algorithm requires the gradient and Hessian of the group-specific log-likelihood $\log p(\mathbf{y}_{i}\mid\bm{\theta})$ at each iteration ${i=1,\dots,N}$, where each $i$ corresponds to a group and $N$ is the total number of groups. In a GLMM setting where latent random effects from different groups are independent, these quantities can be unbiasedly estimated using Fisher’s and Louis’ identities (Cappé et al.,, 2005). However, the latent states in SSMs follow a Markov process and are hence temporally dependent; this dependence makes the computation of the gradient and Hessian of $\log p(\mathbf{y}_{t}\mid\bm{\theta})$, ${t=1,\dots,T}$, challenging. While these quantities can be estimated via particle filtering (e.g., Doucet and Shephard,, 2012), one would need to run a particle filter from the first time period to the current time period for every step in the sequential algorithm. This quickly becomes infeasible for long time series.
Algorithm 1 R-VGAL
Input: exchangeable observations $\mathbf{y}_{1},\dots,\mathbf{y}_{N}$, initial values $\bm{\mu}_{0}$ and $\mathbf{\Sigma}_{0}$.
Output: variational parameters $\bm{\mu}_{i}$ and $\mathbf{\Sigma}_{i}$, for $i=1,...,N$.
Set $q_{0}(\bm{\theta})=N(\bm{\mu}_{0},\mathbf{\Sigma}_{0})$.
for $i=1,\dots,N$ do
[…]
end for
[…]
$l_{W_{k}}(\bm{\theta})=\log f(\omega_{k};\bm{\theta})+\frac{I(\omega_{k})}{f(\omega_{k};\bm{\theta})},\quad k=1,\dots,\left\lfloor\frac{T-1}{2}\right\rfloor,$ (11)
while in the multivariate case,
$l_{W_{k}}(\bm{\theta})=\log\lvert\mathbf{f}(\omega_{k};\bm{\theta})\rvert+\operatorname{Tr}(\mathbf{f}(\omega_{k};\bm{\theta}))^{-1}\mathbf{I}(\omega_{k}),\quad k=1,\dots,\left\lfloor\frac{T-1}{2}\right\rfloor,$ (12)
for $\omega_{k}=\frac{2\pi k}{T}$. The log-likelihood contributions $l_{W_{k}}(\bm{\theta})$ are available for models where the power spectral density is available analytically, such as models where the latent states follow a VARMA (She et al.,, 2022) or a VARTFIMA (Villani et al.,, 2022) process. The gradient ${\nabla_{\bm{\theta}}l_{W_{k}}(\bm{\theta})}$ and Hessian ${\nabla_{\bm{\theta}}^{2}l_{W_{k}}(\bm{\theta})}$ are therefore also available in closed form for these models. At each VB update, ${\nabla_{\bm{\theta}}l_{W_{k}}(\bm{\theta})}$ and ${\nabla_{\bm{\theta}}^{2}l_{W_{k}}(\bm{\theta})}$ are used to update the variational parameters for $k=1,\dots,\left\lfloor\frac{T-1}{2}\right\rfloor$. Their expectations with respect to $q_{k-1}(\bm{\theta})$ are approximated using Monte Carlo samples ${\bm{\theta}^{(s)}\sim q_{k-1}(\bm{\theta})}$, ${s=1,\dots,S}$:
$\mathbb{E}_{q_{k-1}}(\nabla_{\bm{\theta}}l_{W_{k}}(\bm{\theta}))\approx\frac{1}{S}\sum_{s=1}^{S}\nabla_{\bm{\theta}}l_{W_{k}}(\bm{\theta}^{(s)}),\text{ and }\mathbb{E}_{q_{k-1}}(\nabla_{\bm{\theta}}^{2}l_{W_{k}}(\bm{\theta}))\approx\frac{1}{S}\sum_{s=1}^{S}\nabla_{\bm{\theta}}^{2}l_{W_{k}}(\bm{\theta}^{(s)}),$
for $k=1,\dots,\left\lfloor\frac{T-1}{2}\right\rfloor$.
We also implement the damping approach of Vu et al., (2024) to mitigate instability that may arise when traversing the first few frequencies. Damping is done by splitting an update into $D$ steps, where $D$ is user-specific. In each step, the gradient and the Hessian of the log of the Whittle likelihood are multiplied by a “step size” $a=\frac{1}{D}$, and the variational parameters are updated $D$ times during one VB iteration. This has the effect of splitting a log-likelihood contribution into $D$ “parts”, and using them to update variational parameters one part at a time. We summarise R-VGA-Whittle in Algorithm 2.
Algorithm 2 Damped R-VGA-Whittle
Input: time series observations $\mathbf{y}_{1},\dots,\mathbf{y}_{T}$, initial values $\bm{\mu}_{0}$ and $\mathbf{\Sigma}_{0}$, number of frequencies to damp $n_{damp}$, number of damping steps $D$.
Compute the periodogram $\mathbf{I}(\omega_{k})$, $k=-\left\lceil\frac{T}{2}\right\rceil+1,\dots,\left\lfloor\frac{T}{2}\right\rfloor$, via the FFT algorithm.
Set $q_{0}(\bm{\theta})=N(\bm{\mu}_{0},\mathbf{\Sigma}_{0})$.
for $k=1,\dots,\lfloor\frac{T-1}{2}\rfloor$ do
     if $k\leq n_{damp}$ then
         Set $a=1/D$, $\bm{\mu}_{k,0}=\bm{\mu}_{k-1},\mathbf{\Sigma}_{k,0}=\mathbf{\Sigma}_{k-1}$
         for $l=1,\dots,D$ do
              Set $q_{k,l-1}(\bm{\theta})=N(\bm{\mu}_{k,l-1},\mathbf{\Sigma}_{k,l-1})$.
              $\bm{\mu}_{k,l}=\bm{\mu}_{k,l-1}+a\mathbf{\Sigma}_{k,l}\mathbb{E}_{q_{k,l-1}}(\nabla_{\bm{\theta}}l_{W_{k}}(\bm{\theta}))$
              $\mathbf{\Sigma}_{k,l}^{-1}=\mathbf{\Sigma}_{k,l-1}^{-1}-a\mathbb{E}_{q_{k,l-1}}(\nabla_{\bm{\theta}}^{2}l_{W_{k}}(\bm{\theta}))$
         end for
         Set $\bm{\mu}_{k}=\bm{\mu}_{k,D},\mathbf{\Sigma}_{k}=\mathbf{\Sigma}_{k,D}$, $q_{k}(\bm{\theta})=N(\bm{\mu}_{k},\mathbf{\Sigma}_{k})$.
     else
         $\bm{\mu}_{k}=\bm{\mu}_{k-1}+\mathbf{\Sigma}_{k}\mathbb{E}_{q_{k-1}}({\nabla_{\bm{\theta}}l_{W_{k}}(\bm{\theta})})$
         $\mathbf{\Sigma}_{k}^{-1}=\mathbf{\Sigma}_{k-1}^{-1}-\mathbb{E}_{q_{k-1}}({\nabla_{\bm{\theta}}^{2}l_{W_{k}}(\bm{\theta})})$
     end if
end for
Output: variational parameters $\bm{\mu}_{k}$ and $\mathbf{\Sigma}_{k}$, for $k=1,\dots,\lfloor\frac{T-1}{2}\rfloor$.
2.4 R-VGA-Whittle with block updates
The spectral density of a process captures the distribution of signal power (variance) across frequency (Percival and Walden,, 1993). The periodogram is an estimate of the spectral density. In Figure 1(a), we show the periodogram for data generated from the linear Gaussian SSM that we consider in Section 3.1. The variance (power) of the signal gradually decreases as the frequency increases, which means that lower frequencies contribute more to the total power of the process. A similar pattern is seen in Figure 1(b), which shows the periodogram for data generated from a univariate stochastic volatility model that we consider in Section 3.2. In this case there is an even larger concentration of power at the lower frequencies.
(a) Linear Gaussian model.
(b) Univariate SV model.
Figure 1: Plot of raw periodograms of the data (black) and smoothed periodograms (light red) for two univariate models specified in Section 3. Vertical red dashed lines show the frequency cutoffs at half-power (3dB). Frequencies after this cutoff are processed in blocks.
(a) Variable 1.
(b) Variable 2.
Figure 2: Plot of raw periodograms (black) and smoothed periodograms (light red) for the bivariate stochastic volatility model specified in Section 3. Vertical dashed lines show the 3dB frequency cutoff points for each series. In this bivariate case, the cutoff point we use for R-VGA-Whittle is the higher of the two.
Frequencies that carry less power have less effect on R-VGA-Whittle updates. To speed up the algorithm and reduce the number of gradient/Hessian calculations, we therefore process angular frequency contributions with low power, namely those beyond the 3dB cutoff point, together in “blocks”. We define the 3dB cutoff point as the angular frequency at which the power drops to half the maximum power in the (smoothed) periodogram, which we obtain using Welch’s method (Welch,, 1967). Section S3 studies several cutoff points and empirically justifies our choice of a 3dB cutoff.
Assume the 3dB cutoff point occurs at $\omega_{\tilde{n}}$; we split the remaining $\lfloor\frac{T-1}{2}\rfloor-\tilde{n}$ angular frequencies into blocks of approximately equal length $B$. Suppose that there are $n_{b}$ blocks, with $\Omega_{b},b=1,\dots,n_{b}$, denoting the set of angular frequencies in the $b$th block. The variational mean and precision matrix are then updated based on each block’s Whittle log-likelihood contribution. For $b=1,\dots,n_{b}$, this is given by
$l_{W}^{b}(\bm{\theta})=\sum_{k\in\Omega_{b}}l_{W_{k}}(\bm{\theta}),$ (13)
that is, by the sum of the Whittle log-likelihood contributions of the frequencies in the block.
For a multivariate time series, we compute the periodogram for each individual time series along with its corresponding 3dB cutoff point, and then use the highest cutoff point in R-VGA-Whittle. An example of bivariate time series periodograms along with their 3dB cutoffs points is shown in Figure 2.
We summarise R-VGA-Whittle with block updates in Algorithm 3.
Algorithm 3 Block updates for R-VGA with the Whittle likelihood
Input: observations $\mathbf{y}_{1},\dots,\mathbf{y}_{T}$, initial values $\bm{\mu}_{0}$ and $\mathbf{\Sigma}_{0}$, number of angular frequencies to damp $n_{damp}$, number of damping steps $D$, number of angular frequencies to update individually $\tilde{n}>n_{damp}$, block length $B$.
Step 1: Compute the periodogram $\mathbf{I}(\omega_{k})$, $k=-\left\lceil\frac{T}{2}\right\rceil+1,\dots,\left\lfloor\frac{T}{2}\right\rfloor$, via the FFT algorithm.
Step 2: Divide the last $\lfloor\frac{T-1}{2}\rfloor-\tilde{n}$ frequencies into $n_{b}$ blocks of length (approximately) $B$.
Step 3: Set $q_{0}(\bm{\theta})=N(\bm{\mu}_{0},\mathbf{\Sigma}_{0})$.
for $\tilde{k}=1,\dots,\tilde{n}$ do
     Update $\bm{\mu}_{\tilde{k}},\mathbf{\Sigma}_{\tilde{k}}$ as in Algorithm 2
end for
for $b=1,\dots,n_{b}$ do
     Set $\tilde{k}=\tilde{n}+b$
     $\bm{\mu}_{\tilde{k}}=\bm{\mu}_{{\tilde{k}}-1}+\mathbf{\Sigma}_{\tilde{k}}\mathbb{E}_{q_{{\tilde{k}}-1}}(\nabla_{\bm{\theta}}l_{W}^{b}(\bm{\theta}))$, where $l_{W}^{b}(\bm{\theta})$ is given in (13)
     $\mathbf{\Sigma}_{\tilde{k}}^{-1}=\mathbf{\Sigma}_{{\tilde{k}}-1}^{-1}-\mathbb{E}_{q_{{\tilde{k}}-1}}(\nabla_{\bm{\theta}}^{2}l_{W}^{b}(\bm{\theta}))$
end for
Output: variational parameters $\bm{\mu}_{\tilde{k}}$ and $\mathbf{\Sigma}_{\tilde{k}}$, for $\tilde{k}=1,\dots,\tilde{n}+n_{b}$.
2.5 HMC-Whittle
We now discuss the use of HMC with the Whittle likelihood for estimating parameters in SSMs. Suppose we want to sample from the posterior distribution $p(\bm{\theta}\mid\mathbf{y}_{1:T})$. We introduce an auxiliary momentum variable $\mathbf{r}$ of the same length as $\bm{\theta}$ and assume $\mathbf{r}$ comes from a density $N(\mathbf{0},\mathbf{M})$, where $\mathbf{M}$ is a mass matrix that is often a multiple of the identity matrix. The joint density of $\bm{\theta}$ and $\mathbf{r}$ is $p(\bm{\theta},\mathbf{r}\mid\mathbf{y})\propto\exp(-H(\bm{\theta},\mathbf{r}))$, where
$H(\bm{\theta},\mathbf{r})=-\mathcal{L}(\bm{\theta})+\frac{1}{2}\mathbf{r}^{\top}\mathbf{M}^{-1}\mathbf{r}$ (14)
is called the Hamiltonian, and $\mathcal{L}(\bm{\theta})=\log p(\bm{\theta}\mid\mathbf{y}_{1:T})$. HMC proceeds by moving $\bm{\theta}$ and $\mathbf{r}$ according to the Hamiltonian equations
$\frac{\,\mathrm{d}\bm{\theta}}{\,\mathrm{d}t}=\frac{\partial H}{\partial\mathbf{r}}=\mathbf{M}^{-1}\mathbf{r},\quad\frac{\,\mathrm{d}\mathbf{r}}{\,\mathrm{d}t}=-\frac{\partial H}{\partial\bm{\theta}}=\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta}),$ (15)
to obtain new proposed values $\bm{\theta}^{*}$ and $\mathbf{r}^{*}$. By Bayes’ theorem, $p(\bm{\theta}\mid\mathbf{y}_{1:T})\propto p(\mathbf{y}_{1:T}\mid\bm{\theta})p(\bm{\theta})$, so the gradient $\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta})$ can be written as
$\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta})=\nabla_{\bm{\theta}}\log p(\mathbf{y}_{1:T}\mid\bm{\theta})+\nabla_{\bm{\theta}}\log p(\bm{\theta}).$ (16)
The gradient $\nabla_{\bm{\theta}}\log p(\mathbf{y}_{1:T}\mid\bm{\theta})$ on the right hand side of (16) is typically intractable in SSMs and needs to be estimated using a particle filter (e.g., Nemeth et al.,, 2016). However, if the exact likelihood is replaced with the Whittle likelihood, an approximation to ${\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta})}$ can be obtained without the need for running an expensive particle filter repeatedly. That is, rather than targeting the true posterior distribution $p(\bm{\theta}\mid\mathbf{y}_{1:T})$, we target an approximation $\tilde{p}(\bm{\theta}\mid\mathbf{y}_{1:T})\propto p_{W}(\mathbf{y}_{1:T}\mid\bm{\theta})p(\bm{\theta})$, where $p_{W}(\mathbf{y}_{1:T}\mid\bm{\theta})\equiv\exp(l_{W}(\bm{\theta}))$ is the Whittle likelihood as defined in (6) for a univariate time series, or (10) for multivariate time series. Then $\mathcal{L}(\bm{\theta})$ in (14) is replaced by the quantity $\mathcal{L}_{W}(\bm{\theta})=\log\tilde{p}(\bm{\theta}\mid\mathbf{y}_{1:T})$, and the term $\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta})$ in (15) is replaced with
$\nabla_{\bm{\theta}}\mathcal{L}_{W}(\bm{\theta})=\nabla_{\bm{\theta}}l_{W}(\bm{\theta})+\nabla_{\bm{\theta}}\log p(\bm{\theta}).$ (17)
For models where the spectral density is available in closed form, the Whittle likelihood is also available in closed form, and the gradient $\nabla_{\bm{\theta}}\mathcal{L}_{W}(\bm{\theta})$ can be obtained analytically.
In practice, the continuous-time Hamiltonian dynamics in (15) are simulated over $L$ discrete time steps of size $\epsilon$ using a leapfrog integrator (see Algorithm 4). The resulting $\bm{\theta}^{*}$ and $\mathbf{r}^{*}$ from this integrator are then used as proposals and are accepted with probability
$\min\left(1,\frac{\exp(-H(\bm{\theta}^{*},\mathbf{r}^{*})}{\exp(-H(\bm{\theta},\mathbf{r})}\right).$
This process of proposing and accepting/rejecting is repeated until the desired number of posterior samples has been obtained. We summarise HMC-Whittle in Algorithm 5.
Algorithm 4 One step of the Leapfrog algorithm
Input: parameters $\bm{\theta}$, $\mathbf{r}$, leapfrog step size $\epsilon$.
Set $\mathbf{r}^{*}=\mathbf{r}+\epsilon\nabla_{\bm{\theta}}\mathcal{L}\left({\bm{\theta}}\right)/2$,
Set $\bm{\theta}^{*}={\bm{\theta}}+\epsilon\bm{{M}}^{-1}\mathbf{r}^{*}$,
Set $\mathbf{r}^{*}=\mathbf{r}^{*}+\epsilon\nabla_{\bm{\theta}}\mathcal{L}\left(\bm{\theta}^{*}\right)/2$.
return $\bm{\theta}^{*}$, $\mathbf{r}^{*}$.
Algorithm 5 HMC-Whittle
Input: observations $\mathbf{y}_{1},\dots,\mathbf{y}_{T}$, prior distribution $p(\bm{\theta})$, mass matrix $\mathbf{M}$, number of HMC iterations $J$, leapfrog step size $\epsilon$, number of leapfrog steps $L$.
Sample $\bm{\theta}^{(0)}\sim p(\bm{\theta})$.
for $j=1,\dots,J$ do
     Sample $\mathbf{r}^{(0)}\sim N(\mathbf{0},\mathbf{M})$.
     Set $\bm{\theta}^{*}=\bm{\theta}^{(j-1)}$, $\mathbf{r}^{*}=\mathbf{r}^{(0)}$.
     for $l=1,...,L$ do
         $(\bm{\theta}^{*},\mathbf{r}^{*})\leftarrow\text{Leapfrog}(\bm{\theta}^{*},\mathbf{r}^{*},\epsilon)$. (see Algorithm 4)
     end for
     With probability
$\alpha=\text{min}\left(1,\frac{\exp\left(\mathcal{L}_{W}(\bm{\theta}^{*})-\frac{1}{2}\mathbf{r}^{*\top}\mathbf{M}^{-1}\mathbf{r}^{*}\right)}{\exp\left(\mathcal{L}_{W}(\bm{\theta}^{(j-1)})-\frac{1}{2}\mathbf{r}^{(0)\top}\mathbf{M}^{-1}\mathbf{r}^{(0)}\right)}\right),$
     set $\bm{\theta}^{(j)}=\bm{\theta}^{*}$, else $\bm{\theta}^{(j)}=\bm{\theta}^{(j-1)}$.
end for
Output: posterior samples $\bm{\theta}^{(1)},\dots,\bm{\theta}^{(J)}$.
3 Applications
This section demonstrates the use of R-VGA-Whittle for estimating parameters of a univariate linear Gaussian SSM and those of a univariate/bivariate stochastic volatility (SV) model using simulated and real data. We compare the posterior densities obtained from R-VGA-Whittle to those from HMC-exact and HMC-Whittle. All examples are implemented in R, and the reproducible code is available from https://github.com/bao-anh-vu/R-VGA-Whittle. HMC-exact and HMC-Whittle are implemented using the CmdStanR interface to the Stan programming language (Gabry et al.,, 2024).
For all examples, we implement R-VGA-Whittle with damping and blocking as described in Algorithm 3. We apply damping to the first $n_{damp}=5$ angular frequencies with $D=100$ damping steps each. These values were chosen to reduce the initial instability at the expense of a small additional computational cost. We select $\tilde{n}$ for each model by finding the 3dB frequency cutoff and process the remaining frequencies in blocks of length $B=100$, following a study in Section S3 of the online supplement. At each iteration of R-VGA-Whittle, we use $S=1000$ Monte Carlo samples to approximate the expectations of the gradient and Hessian of the Whittle log-likelihood. This value of $S$ was chosen because it results in very little variation between repeated runs of R-VGA-Whittle; see Section S4 of the online supplement for an experiment exploring the algorithm’s sensitivity to $S$. HMC-Whittle and HMC-exact were run with two chains with 15000 samples each. The first 5000 samples of each chain were discarded as burn-in. All experiments were carried out on a high-end server, with an NVIDIA Tesla V100 32GB graphics processing unit (GPU) used to parallelise over the $S=1000$ Monte Carlo samples needed for estimating the expectations of the gradient and Hessian of the Whittle log-likelihood at each iteration of Algorithm 2.
3.1 Univariate linear Gaussian SSM
We generate $T=10000$ observations from the linear Gaussian SSM,
$y_{t}$ $=x_{t}+\epsilon_{t},\quad\epsilon_{t}\sim N(0,\sigma_{\epsilon}^{2}),\quad t=1,\dots,T,$ (18)
$x_{t}$ $=\phi x_{t-1}+\eta_{t},\quad\eta_{t}\sim N(0,\sigma_{\eta}^{2}),\quad t=2,\dots,T,$ (19)
$x_{1}$ $\sim N\left(0,\frac{\sigma_{\eta}^{2}}{1-\phi^{2}}\right).$ (20)
We set the true auto-correlation parameter to $\phi=0.9$, the true state disturbance standard deviation to $\sigma_{\eta}=0.7$ and the true measurement error standard deviation to $\sigma_{\epsilon}=0.5$.
The parameters that need to be estimated, $\phi,\sigma_{\eta}$ and $\sigma_{\epsilon}$, take values in subsets of $\mathbb{R}$. We therefore instead estimate the transformed parameters ${\bm{\theta}=(\theta_{\phi},\theta_{\eta},\theta_{\epsilon})^{\top}}$, where $\theta_{\phi}=\tanh^{-1}(\phi)$, $\theta_{\eta}=\log(\sigma_{\eta}^{2})$, and $\theta_{\epsilon}=\log(\sigma_{\epsilon}^{2})$, which are unconstrained.
The spectral density of the observed time series $\{y_{t}\}_{t=1}^{T}$ is
$f_{y}(\omega_{k};\bm{\theta})$ $=f_{x}(\omega_{k};\bm{\theta})+f_{\epsilon}(\omega_{k};\bm{\theta}),\quad\omega_{k}=\frac{2\pi k}{T},\quad k=-\left\lceil\frac{T}{2}\right\rceil+1,\dots,\left\lfloor\frac{T}{2}\right\rfloor,$ (21)
where
$f_{x}(\omega_{k};\bm{\theta})=\frac{\sigma_{\eta}^{2}}{1+\phi^{2}-2\phi\cos(\omega_{k})}$
is the spectral density of the AR(1) process $\{x_{t}\}_{t=1}^{T}$, and
$f_{\epsilon}(\omega_{k};\bm{\theta})=\sigma_{\epsilon}^{2}$
is the spectral density of the white noise process $\{\epsilon_{t}\}_{t=1}^{T}$. The Whittle log-likelihood for this model is given by (6), where the periodogram $\{I(\omega_{k}):k=-\left\lceil\frac{T}{2}\right\rceil+1,\dots,\left\lfloor\frac{T}{2}\right\rfloor\}$ is the discrete Fourier transform (DFT) of the observations $\{y_{t}\}_{t=1}^{T}$, which we compute using the fft function in R. We evaluate the gradient and Hessian of the Whittle log-likelihood at each frequency, $\nabla_{\theta}l_{W_{k}}(\bm{\theta})$ and $\nabla_{\theta}^{2}l_{W_{k}}(\bm{\theta})$, using automatic differentiation via the tensorflow library (Abadi et al.,, 2015) in R. The initial/prior distribution we use is
$p(\bm{\theta})=q_{0}(\bm{\theta})=N\left(\begin{bmatrix}0\\ -1\\ -1\end{bmatrix},\begin{bmatrix}1&0&0\\ 0&1&0\\ 0&0&1\\ \end{bmatrix}\right).$ (22)
This leads to 95% probability intervals of $(-0.96,0.96)$ for $\phi$, and $(0.23,1.59)$ for $\sigma_{\eta}$ and $\sigma_{\epsilon}$.
Figure 3 shows the trajectory of the R-VGA-Whittle variational mean for each parameter. There are only 226 R-VGA-Whittle updates, as the 3dB cutoff point is the 176th angular frequency, and there are 50 blocks thereafter. The figure shows that the trajectory of $\phi$ moves towards the true value within the first 10 updates. This is expected, since most of the information on $\phi$, which is reasonably large in our example, is in the low-frequency components of the signal. The trajectories of $\sigma_{\eta}$ and $\sigma_{\epsilon}$ are more sporadic before settling close to the true value around the last 25 (block) updates. This later convergence is not unexpected since high-frequency log-likelihood components contain non-negligible information on these parameters that describe fine components of variation (the spectral density of white noise is a constant function of frequency). This behaviour is especially apparent for the measurement error standard deviation $\sigma_{\epsilon}$.
Figures 8(a) and 8(b) in Section S5 of the online supplement show the trace plots for each parameter obtained from HMC-exact and HMC-Whittle. Trace plots from both methods show good mixing for all three parameters but, for $\sigma_{\eta}$ and $\sigma_{\epsilon}$ in particular, HMC-Whittle has slightly better mixing than HMC-exact.
Figure 4 shows the marginal posterior distributions of the parameters, along with bivariate posterior distributions. The three methods yield very similar posterior distributions for all parameters. We also show in Section S3 of the online supplement that the posterior densities obtained from R-VGA-Whittle using block size $B=100$ are highly similar to those obtained without block updates. This suggests that block updates are useful for speeding up R-VGA-Whittle without compromising the accuracy of the approximate posterior densities.
Figure 3: Trajectories of the variational means for parameters in the linear Gaussian SSM with simulated data. True parameter values are marked with dashed lines. The vertical line is at $\tilde{n}$, the point beyond which updates are done in blocks.
Figure 4: Posterior distributions from HMC-exact (blue), and approximate posterior distributions from R-VGA-Whittle (red) and HMC-Whittle (yellow) for data simulated from the univariate linear Gaussian SSM with $T=10000$. Diagonal panels: Marginal posterior distributions with true parameters denoted using dotted lines. Off-diagonal panels: Bivariate posterior distributions with true parameters denoted using the symbol $\times$.
$T$ R-VGA-Whittle HMC-Whittle HMC-exact
Linear Gaussian 10000 14 125 447
SV (sim. data) 10000 13 104 17005
SV (real data) 2649 13 25 3449
2D SV (sim. data) 5000 121 14521 72170
2D SV (real data) 2649 107 8890 74462
Table 1: Computing time (in seconds) for R-VGA-Whittle, HMC-Whittle and HMC-exact for each model we considered. The number of observations $T$ for each example is also shown.
A comparison of the computing time between R-VGA-Whittle, HMC-Whittle and HMC-exact for this model can be found in Table 1. The table shows that R-VGA-Whittle is nearly 10 times faster than HMC-Whittle and more than 30 times faster than HMC-exact.
3.2 Univariate stochastic volatility model
SV models are a popular class of non-Gaussian financial time series models used for modelling the volatility of stock returns. An overview of SV models and traditional likelihood-based methods for parameter inference for use with these models can be found in Kim et al., (1998); an overview of MCMC-based inference for a univariate SV model can be found in Fearnhead, (2011).
In this section, we follow Fearnhead, (2011) and simulate $T=10000$ observations from the following univariate SV model:
$y_{t}$ $=\sigma_{t}\epsilon_{t},\quad\epsilon_{t}\sim N(0,1),\quad t=1,\dots,T,$ (23)
$\sigma_{t}$ $=\kappa\exp\left(\frac{x_{t}}{2}\right),\quad t=1,\dots,T,$ (24)
$x_{t}$ $=\phi x_{t-1}+\eta_{t},\quad\eta_{t}\sim N(0,\sigma_{\eta}^{2}),\quad t=2,\dots,T,$ (25)
$x_{1}$ $\sim N\left(0,\frac{\sigma_{\eta}^{2}}{1-\phi^{2}}\right).$ (26)
We set the true parameters to $\phi=0.99$, $\sigma_{\eta}=0.1$, and $\kappa=2$.
This model can be “linearised” with a log-squared transformation (Ruiz,, 1994),
$\log(y_{t}^{2})$ $=\log(\kappa^{2})+x_{t}+\log(\epsilon_{t}^{2})$
$=\log(\kappa^{2})+\mathbb{E}(\log(\epsilon_{t}^{2}))+x_{t}+\log(\epsilon_{t}^{2})-\mathbb{E}(\log(\epsilon_{t}^{2}))$
$=\mu+x_{t}+\xi_{t},$ (27)
where $\mu\equiv\log(\kappa^{2})+\mathbb{E}(\log(\epsilon_{t}^{2}))$ and $\xi_{t}\equiv\log(\epsilon_{t}^{2})-\mathbb{E}(\log(\epsilon_{t}^{2}))$, for $t=1,\dots,T$. The terms $\{\xi_{t}\}_{t=1}^{T}$ are independent and identically distributed (iid) non-Gaussian terms with mean zero and variance $\frac{\pi^{2}}{2}$ (see Ruiz, (1994) for a derivation, which is reproduced in Section S2 of the online supplement for completeness).
Rearranging (27) yields the de-meaned series
$z_{t}\equiv\log(y_{t}^{2})-\mu=x_{t}+\xi_{t},\quad t=1,\dots,T.$ (28)
Since $\mathbb{E}(\log(y_{t}^{2}))=\mu$, the parameter $\mu$ can be estimated using the sample mean:
$\hat{\mu}=\frac{1}{T}\sum_{t=1}^{T}\log(y_{t}^{2}).$
As $\mu=\log(\kappa^{2})+\mathbb{E}(\log(\epsilon_{t}^{2}))$, once an estimate $\hat{\mu}$ is obtained, $\kappa$ can be estimated as
$\hat{\kappa}=\sqrt{\exp(\hat{\mu}-\mathbb{E}(\log(\epsilon_{t}^{2})))},$
where $\mathbb{E}(\log(\epsilon_{t}^{2})))=\psi(1/2)+\log(2)$, and $\psi(\cdot)$ is the digamma function (Abramowitz and Stegun,, 1968). As we show below, the estimate $\hat{\kappa}$ is not required for R-VGA-Whittle and HMC-Whittle as the Whittle likelihood does not depend on $\kappa$. However, $\kappa$ appears in the likelihood used by HMC-exact. Since $\kappa$ is not estimated by the Whittle-based methods, for comparison purposes we fix it to its estimate $\hat{\kappa}$ when using HMC-exact.
The remaining parameters to estimate are $\phi$ and $\sigma_{\eta}$. As with the linear Gaussian state space model, we work with the transformed parameters $\bm{\theta}=(\theta_{\phi},\theta_{\eta})^{\top}$, where $\theta_{\phi}=\tanh^{-1}(\phi)$ and $\theta_{\eta}=\log(\sigma_{\eta}^{2})$, which are unconstrained. We fit the spectral density
$f_{z}(\omega_{k};\bm{\theta})$ $=f_{x}(\omega_{k};\bm{\theta})+f_{\xi}(\omega_{k};\bm{\theta}),$ (29)
to the log-squared, demeaned series $\{z_{t}\}_{t=1}^{T}$, where
$f_{x}(\omega_{k};\bm{\theta})=\frac{\sigma_{\eta}^{2}}{1+\phi^{2}-2\phi\cos(\omega_{k})},$
and $f_{\xi}(\omega_{k};\bm{\theta})$ is the Fourier transform of the covariance function $C_{\xi}(h;\bm{\theta})$ of $\{\xi_{t}\}_{t=1}^{T}$:
$C_{\xi}(h;\bm{\theta})=\mathbb{E}(\xi_{t}\xi_{t-h})=\begin{cases}\frac{\pi^{2}}{2},&h=0,\\ 0,&h\neq 0,\end{cases}$
and is thus given by
$f_{\xi}(\omega_{k};\bm{\theta})=\sum_{h=-\infty}^{\infty}C_{\xi}(h;\bm{\theta})\exp(-i\omega_{k}h)=C_{\xi}(0)=\frac{\pi^{2}}{2},\quad k=-\left\lceil\frac{T}{2}\right\rceil+1,\dots,\left\lfloor\frac{T}{2}\right\rfloor.$ (30)
The prior/initial variational distribution we use is
$p(\bm{\theta})=q_{0}(\bm{\theta})=N\left(\begin{bmatrix}2\\ -3\end{bmatrix},\begin{bmatrix}0.5&0\\ 0&0.5\\ \end{bmatrix}\right),$ (31)
which leads to 95% probability intervals of (0.77, 0.99) for $\phi$, and (0.14, 0.36) for $\sigma_{\eta}$.
Figure 5 shows the trajectories of the R-VGA-Whittle variational means for all parameters. The 3dB cutoff point for this example is the 75th frequency, after which there are 50 blocks, for a total of 125 updates. Similar to the case of the LGSS model in Section 3.1, the variational mean of $\phi$ is close to the true value from the early stages of the algorithm, while that of $\sigma_{\eta}$ requires more iterations to achieve convergence.
Figures 9(a)9(c) in Section S5 of the online supplement show the HMC-exact and HMC-Whittle trace plots. For this example, HMC-exact produces posterior samples with a much higher correlation than HMC-Whittle, particularly for the variance parameter $\sigma_{\eta}$. This is because HMC-Whittle only targets the posterior density of the parameters, while HMC-exact targets the joint posterior density of the states and parameters, which are high dimensional. We thin the posterior samples from HMC-exact by a factor of 100 for the remaining samples to be treated as uncorrelated draws from the posterior distribution.
We compare the posterior densities from R-VGA-Whittle, HMC-Whittle, and HMC-exact in Figure 6. The posterior densities from R-VGA-Whittle and HMC-Whittle are very similar, while the posterior densities from HMC-exact are slightly narrower. Table 1 shows that R-VGA-Whittle is 8 times faster than HMC-Whittle, and more than 1000 times faster than HMC-exact in this example.
Figure 5: Trajectories of the variational means for parameters in the univariate SV model with simulated data. True parameter values are marked with dashed lines. The vertical line is at $\tilde{n}$, the point beyond which updates are done in blocks.
Figure 6: Posterior distributions from HMC-exact (blue), and approximate posterior distributions from R-VGAL-Whittle (red) and HMC-Whittle (yellow) for data simulated from the univariate SV model with $T=10000$. Diagonal panels: Marginal posterior distributions with true parameters denoted using dotted lines. Off-diagonal panels: Bivariate posterior distributions with true parameters denoted using the symbol $\times$.
3.3 Bivariate SV model
We now consider the following bivariate extension of the SV model in Section 3.2 and simulate $T=5000$ observations as follows:
$\mathbf{y}_{t}$ $=\mathbf{V}_{t}\bm{\varepsilon}_{t},\quad\bm{\varepsilon}_{t}\sim N(\mathbf{0},\mathbf{I}_{2}),\quad t=1,\dots,T,$ (32)
$\mathbf{x}_{t}$ $=\mathbf{\Phi}\mathbf{x}_{t-1}+\bm{\eta}_{t},\quad\bm{\eta}_{t}\sim N(\mathbf{0},\mathbf{\Sigma}_{\eta}),\quad t=2,\dots,T,$ (33)
$\mathbf{x}_{1}$ $\sim N(\mathbf{0},\mathbf{\Sigma}_{\eta_{1}}),$ (34)
where $\mathbf{x}_{t}\equiv(x_{1,t},x_{2,t})^{\top}$, $\mathbf{y}_{t}\equiv(y_{1,t},y_{2,t})^{\top}$, ${\mathbf{V}_{t}\equiv\operatorname{diag}(\exp(\tfrac{x_{1,t}}{2}),\exp(\tfrac{x_{2,t}}{2}))}$, and $\mathbf{I}_{m}$ is the $m\times m$ identity matrix. The covariance matrix $\mathbf{\Sigma}_{\eta_{1}}$ is the solution to the equation $\operatorname{vec}(\mathbf{\Sigma}_{\eta_{1}})=(\mathbf{I}_{4}-\mathbf{\Phi}\otimes\mathbf{\Phi})^{-1}\operatorname{vec}(\mathbf{\Sigma}_{\eta})$, where $\textrm{vec}(\cdot)$ is the vectorisation operator and $\otimes$ denotes the Kronecker product of two matrices. We let $\mathbf{\Phi}$ be a diagonal matrix but let $\mathbf{\Sigma}_{\eta}$ have off-diagonal entries so that there is dependence between the two series. We set these matrices to
$\mathbf{\Phi}=\begin{bmatrix}0.99&0\\ 0&0.98\end{bmatrix},\quad\mathbf{\Sigma}_{\eta}=\begin{bmatrix}0.02&0.005\\ 0.005&0.01\end{bmatrix}.$ (35)
As with the univariate model, we take a log-squared transformation of (32) to find
$\tilde{\mathbf{y}}_{t}=\bm{\mu}+\mathbf{x}_{t}+\bm{\xi}_{t},$ (36)
where
$\tilde{\mathbf{y}}_{t}=\begin{bmatrix}\log(y_{1,t}^{2})\\ \log(y_{2,t}^{2})\end{bmatrix},\quad\bm{\mu}=\begin{bmatrix}\mu_{1}\\ \mu_{2}\end{bmatrix}=\begin{bmatrix}\mathbb{E}(\log(\epsilon_{1,t}^{2}))\\ \mathbb{E}(\log(\epsilon_{2,t}^{2}))\end{bmatrix},\quad\bm{\xi}_{t}=\begin{bmatrix}\xi_{1,t}\\ \xi_{2,t}\end{bmatrix}=\begin{bmatrix}\log(\epsilon_{1,t}^{2})-\mathbb{E}(\log(\epsilon_{1,t}^{2}))\\ \log(\epsilon_{2,t}^{2})-\mathbb{E}(\log(\epsilon_{2,t}^{2}))\\ \end{bmatrix}.$
The terms $\xi_{1,t}$ and $\xi_{2,t}$ are iid non-Gaussian uncorrelated terms, each with mean zero and variance $\sigma_{\xi}^{2}=\frac{\pi^{2}}{2}$. We express (36) as
$\mathbf{z}_{t}\equiv\tilde{\mathbf{y}}_{t}-\bm{\mu}=\mathbf{x}_{t}+\bm{\xi}_{t},$
and use the sample means
$\hat{\mu}_{1}=\frac{1}{T}\sum_{t=1}^{T}\log(y_{1,t}^{2}),\quad\hat{\mu}_{2}=\frac{1}{T}\sum_{t=1}^{T}\log(y_{2,t}^{2})$
as estimates of $\mu_{1}$ and $\mu_{2}$, respectively. The parameters in this model are $\mathbf{\Phi}$ and $\mathbf{\Sigma}_{\eta}$. To estimate these parameters, we fit the following spectral density to $\{\mathbf{z}_{t}\}_{t=1}^{T}$:
$\mathbf{f}_{\mathbf{z}}(\omega_{k};\bm{\theta})=\mathbf{f}_{\mathbf{x}}(\omega_{k};\bm{\theta})+\mathbf{f}_{\bm{\xi}}(\omega_{k};\bm{\theta}),$
where $\mathbf{f}_{\mathbf{x}}(\omega_{k};\bm{\theta})$ is the spectral density matrix of the VAR(1) process $\{\mathbf{x}_{t}\}_{t=1}^{T}$,
$\mathbf{f}_{\mathbf{x}}(\omega_{k};\bm{\theta})=(\mathbf{I}_{2}-\mathbf{\Phi}e^{-i\omega_{k}})^{-1}\mathbf{\Sigma}_{\eta}(\mathbf{I}_{2}-\mathbf{\Phi}e^{-i\omega_{k}})^{-H},\quad k=-\left\lceil\frac{T}{2}\right\rceil+1,\dots,\left\lfloor\frac{T}{2}\right\rfloor,$ (37)
with $\mathbf{A}^{-H}$ denoting the inverse of the conjugate transpose of the matrix $\mathbf{A}$, and $\mathbf{f}_{\bm{\xi}}(\omega_{k};\bm{\theta})=\tfrac{\pi^{2}}{2}\mathbf{I}_{2}$ is the spectral density of $\{\bm{\xi}_{t}\}_{t=1}^{T}$ (following a similar derivation as that for the univariate SV model in Equation (30)).
We parameterise $\theta_{ii}=\tanh^{-1}(\Phi_{ii}),i=1,2$, ${\mathbf{\Sigma}_{\eta}=\mathbf{L}\mathbf{L}^{\top}}$, where $\mathbf{L}$ is the lower Cholesky factor of $\mathbf{\Sigma}_{\eta}$, and the diagonal elements of $\mathbf{L}$ as $\gamma_{ii}=\log(l_{ii}),i=1,2$. The vector of transformed parameters is ${\bm{\theta}=(\theta_{11},\theta_{22},\gamma_{11},\gamma_{22},l_{21})^{\top}}$.
The initial/prior distribution we use is
$p(\bm{\theta})=q_{0}(\bm{\theta})=N\left(\begin{bmatrix}2\\ 2\\ -2\\ -3\\ 0\end{bmatrix},\begin{bmatrix}0.5&0&0&0&0\\ 0&0.5&0&0&0\\ 0&0&0.5&0&0\\ 0&0&0&0.05&0\\ 0&0&0&0&0.05\\ \end{bmatrix}\right).$ (38)
This leads to 95% probability intervals of $(0.52,0.99)$ for $\Phi_{11}$ and $\Phi_{22}$, $(0.001,0.286)$ for $\Sigma_{\eta_{11}}$, $(-0.101,0.101)$ for $\Sigma_{\eta_{21}}$, and $(0.002,0.261)$ for $\Sigma_{\eta_{22}}$.
Figure 7: Trajectories of the variational means for parameters in the bivariate SV model with simulated data. True parameter values are marked with dashed lines. The vertical line is at $\tilde{n}$, the point beyond which updates are done in blocks.
Figure 7 shows the trajectories of the R-VGA-Whittle variational means for all parameters in this model. For most parameters, the trajectories quickly move towards the true value within the first 10 to 20 updates, with only small adjustments afterwards. The only exception is for parameter $\Sigma_{\eta_{11}}$, which exhibits more gradual changes before settling close to the true value in the last 25 iterations; this is somewhat expected since this parameter corresponds to a high-frequency signal component. R-VGA-Whittle slightly overestimated $\Phi_{22}$ and underestimated $\Sigma_{\eta_{22}}$.
Figures 10(a)10(c) in Section S5 of the online supplement show the HMC-exact and HMC-Whittle trace plots for this model. HMC-exact samples are highly correlated, especially those of $\Sigma_{\eta_{22}}$. We find that it is necessary to thin the HMC-exact samples by a factor of 50 for the samples to be considered uncorrelated draws from the posterior distribution. In contrast, HMC-Whittle produces well-mixed posterior samples. Section S5 of the online supplement also shows that the effective sample sizes from HMC-Whittle are much higher than those from HMC-exact for all parameters.
We compare the posterior densities from R-VGA-Whittle, HMC-Whittle, and HMC-exact in Figure 8. In this example, we observe broad agreement between the three methods, though the posterior densities from R-VGA-Whittle and HMC-Whittle are wider than those from HMC-exact. The posterior density of $\Phi_{22}$ from R-VGA-Whittle puts more probability on values close to one than those from the other two methods, while the posterior density of $\Sigma_{\eta_{21}}$ from both R-VGA-Whittle and HMC-Whittle appear slightly shifted towards zero compared to that from HMC-exact. As shown in Table 1, R-VGA-Whittle is nearly 600 times faster than HMC-exact, and over 100 times faster than HMC-Whittle in this example.
Figure 8: Posterior distributions from HMC-exact (blue), and approximate posterior distributions from R-VGA-Whittle (red) and HMC-Whittle (yellow) for data simulated from the bivariate SV model with $T=5000$. Diagonal panels: Marginal posterior distributions with true parameters denoted using dotted lines. Off-diagonal panels: Bivariate posterior distributions with true parameters denoted using the symbol $\times$.
3.4 Real data examples
We apply R-VGA-Whittle to estimate parameters in univariate and bivariate stochastic volatility models from real data. We first apply it to daily JPY-EUR exchange rate data (Kastner et al.,, 2017) from April 1st, 2005 to August 6th, 2015, which consists of $2650$ observations. Typically, the data used in SV models are de-meaned log-returns, which are computed from the raw exchange rate values as:
$y_{t}=y_{t}^{*}-\frac{1}{T}\sum_{t^{\prime}=1}^{T}y_{t^{\prime}}^{*},\quad y_{t}^{*}=\log\left(\frac{r_{t+1}}{r_{t}}\right),\quad t=1,\dots,2649,$
where $r_{t}$ is the exchange rate on day $t$, and $y_{t}^{*}$ is the log-return on day $t$. We fit the model (23)–(26) to this time series of de-meaned log-returns with $T=2649$, and use the initial variational distribution specified in (31).
The HMC-exact trace plots in Figures 11(a) of the online supplement show that posterior samples of $\phi$ are relatively well-mixed, but that the posterior samples of $\sigma_{\eta}$ are highly correlated. We find it necessary to thin these samples by a factor of 100 in order for them to be considered uncorrelated. In contrast, HMC-Whittle produces well-mixed posterior samples, so that no thinning is needed.
We compare the posterior densities from R-VGA-Whittle, HMC-Whittle and HMC-exact in Figure 9. The posterior densities obtained from R-VGA-Whittle are similar, but slightly wider than those obtained from HMC-Whittle and HMC-exact.
Figure 9: Posterior distributions from HMC-exact (blue), and approximate posterior distributions from R-VGAL-Whittle (red) and HMC-Whittle (yellow) for the univariate SV model fitted using JPY-EUR exchange rate data with $T=2649$. Diagonal panels: Marginal posterior distributions of the parameters. Off-diagonal panels: Bivariate posterior distributions of the parameters.
We also apply R-VGA-Whittle to a bivariate time series of daily GBP-EUR and USD-EUR exchange rates (Kastner et al.,, 2017) from April 1st, 2005, to August 6th, 2015, which consists of $2650$ observations per exchange rate. We compute the de-meaned log-returns for both series and then fit the model (32)–(34) to these log-returns. The prior distribution we use is specified in (38).
Figures 12(a)12(c) in Section S5 of the online supplement show the HMC-exact and HMC-Whittle trace plots for the bivariate SV model with exchange rate data. Similar to the simulation study, we find that HMC-exact samples are highly correlated; samples of parameters in $\mathbf{\Phi}$ tend to be less correlated than samples of parameters in $\mathbf{\Sigma}_{\eta}$. We thin the HMC-exact samples by a factor of 50 so that the remaining samples can be treated as uncorrelated. Posterior samples from HMC-Whittle have low correlation and are mixed well for all parameters, so no thinning is necessary.
We compare the posterior densities from R-VGA-Whittle, HMC-Whittle, and HMC-exact in Figure 10. The marginal posterior densities from R-VGA-Whittle and HMC-Whittle are similar, but noticeably wider than the posterior densities from HMC-exact. Nevertheless, there is a substantial overlap between the posterior densities from all three methods.
Table 1 shows that R-VGA-Whittle is twice as fast as HMC-Whittle in the univariate case, and more than 80 times faster than HMC-Whittle in the bivariate case. R-VGA-Whittle is significantly faster than HMC-exact, with approximately a 260-fold speedup in the univariate case and a 700-fold speedup in the bivariate cases.
Figure 10: Posterior distributions from HMC-exact (blue), and approximate posterior distributions from R-VGAL-Whittle (red) and HMC-Whittle (yellow) for the bivariate SV model fitted using GBP-EUR and USD-EUR exchange rate data with $T=2649$. Diagonal panels: Marginal posterior distributions of the parameters. Off-diagonal panels: Bivariate posterior distributions of the parameters.
4 Conclusion
We develop a novel sequential variational Bayes algorithm that uses the Whittle likelihood, called R-VGA-Whittle, for parameter inference in linear non-Gaussian state space models. The Whittle likelihood is a frequency-domain approximation of the exact likelihood. It is available in closed form and its log can be expressed as a sum of individual log-likelihood frequency components. At each iteration, R-VGA-Whittle sequentially updates the variational parameters by processing one frequency or a block of frequencies at a time. We also demonstrate the use of the Whittle likelihood in HMC (HMC-Whittle), and compare the parameter estimates from R-VGA-Whittle and HMC-Whittle to those of HMC with the exact likelihood (HMC-exact). We show that the posterior densities from the three methods are generally similar, though posterior densities from R-VGA-Whittle and HMC-Whittle tend to be wider compared to those from HMC-exact. As HMC-Whittle only targets the parameters and not the states, it tends to produce posterior samples of the parameters that are less correlated than those from HMC-exact. We also show that R-VGA-Whittle and HMC-Whittle are much more computationally efficient than HMC-exact. R-VGA-Whittle is particularly efficient when dealing with bivariate time series, where it can be up to two orders of magnitude faster than HMC-Whittle and three orders of magnitude faster than HMC-exact.
There are several avenues for future research. First, R-VGA-Whittle assumes a Gaussian variational distribution, so future work could attempt to extend this framework to more flexible variational distributions for approximating skewed or multimodal posterior distributions. Second, some tuning parameters of R-VGA-Whittle (such as the number of damping observations and number of damping steps) are currently user-defined, so an adaptive tuning scheme could be developed to select these parameters. Third, as the Whittle likelihood is known to produce biased parameter estimates for finite sample sizes, using R-VGA-Whittle with the debiased Whittle likelihood of Sykulski et al., (2019) may improve parameter estimates.
Acknowledgements
This work was supported by ARC SRIEAS Grant SR200100005 Securing Antarctica’s Environmental Future. Andrew Zammit-Mangion’s research was also supported by ARC Discovery Early Career Research Award DE180100203.
S1 The R-VGA and R-VGAL algorithms
In this section, we provide a sketch of the derivations for the R-VGA algorithm of Lambert et al., (2022), which serves as the basis for both the R-VGAL algorithm of Vu et al., (2024) and our R-VGA-Whittle algorithm. This section closely follows Section 2.2.2 in Vu et al., (2024).
The R-VGA algorithm of Lambert et al., (2022) requires iid observations, denoted by ${\mathbf{y}_{1:N}\equiv(\mathbf{y}_{1}^{\top},\dots,\mathbf{y}_{N}^{\top})^{\top}}$, and approximates the posterior distribution $p(\bm{\theta}\mid\mathbf{y}_{1:N})$ with a Gaussian distribution $N(\bm{\mu},\mathbf{\Sigma})$. By assumption of conditional independence between observations $\mathbf{y}_{1:N}$ given the parameters $\bm{\theta}$, the KL divergence between the variational distribution $q_{i}(\bm{\theta})$ and the posterior distribution $p(\bm{\theta}\mid\mathbf{y}_{1:i})$ can be expressed as
$\text{KL}(q_{i}(\bm{\theta})\;\|\;p(\bm{\theta}\mid\mathbf{y}_{1:i}))$ $\equiv\int q_{i}(\bm{\theta})\log\frac{q_{i}(\bm{\theta})}{p(\bm{\theta}\mid\mathbf{y}_{1:i})}\,\mathrm{d}\bm{\theta}$
$=\mathbb{E}_{q_{i}}\left(\log q_{i}(\bm{\theta})-\log p(\bm{\theta}\mid\mathbf{y}_{1:i-1})-\log p(\mathbf{y}_{i}\mid\bm{\theta})\right)+\log p(\mathbf{y}_{1:i})-\log p(\mathbf{y}_{1:i-1}).$
In a sequential VB framework, the posterior distribution after incorporating the first $i-1$ observations, $p(\bm{\theta}\mid\mathbf{y}_{1:i-1})$, is approximated by the variational distribution $q_{i-1}(\bm{\theta})$ to give
$\text{KL}(q_{i}(\bm{\theta})\;\|\;p(\bm{\theta}\mid\mathbf{y}_{1:i}))\approx\mathbb{E}_{q_{i}}(\log q_{i}(\bm{\theta})-\log q_{i-1}(\bm{\theta})-\log p(\mathbf{y}_{i}\mid\bm{\theta}))+\log p(\mathbf{y}_{1:i})-\log p(\mathbf{y}_{1:i-1}).$ (S1)
As the last two terms on the right hand side of (S1) do not depend on $\bm{\theta}$, the KL-minimisation problem is equivalent to solving
$\operatorname*{arg\,min}_{\bm{\mu}_{i},\mathbf{\Sigma}_{i}}\,\mathbb{E}_{q_{i}}(\log q_{i}(\bm{\theta})-\log q_{i-1}(\bm{\theta})-\log p(\mathbf{y}_{i}\mid\bm{\theta})).$ (S2)
Differentiating the expectation in (S2) with respect to $\bm{\mu}_{i}$ and $\mathbf{\Sigma}_{i}$, setting the derivatives to zero, and rearranging the resulting equations, yields the following recursive updates for the variational mean $\bm{\mu}_{i}$ and precision matrix $\mathbf{\Sigma}_{i}^{-1}$:
$\bm{\mu}_{i}$ $=\bm{\mu}_{i-1}+\mathbf{\Sigma}_{i-1}\nabla_{\bm{\mu}_{i}}\mathbb{E}_{q_{i}}(\log p(\mathbf{y}_{i}\mid\bm{\theta})),$ (S3)
$\mathbf{\Sigma}_{i}^{-1}$ $=\mathbf{\Sigma}_{i-1}^{-1}-2\nabla_{\mathbf{\Sigma}_{i}}\mathbb{E}_{q_{i}}(\log p(\mathbf{y}_{i}\mid\bm{\theta})).$ (S4)
Then, using Bonnet’s Theorem (Bonnet,, 1964) on (S3) and Price’s Theorem (Price,, 1958) on (S4), we rewrite the gradient terms as
$\nabla_{\bm{\mu}_{i}}\mathbb{E}_{q_{i}}(\log p(\mathbf{y}_{i}\mid\bm{\theta}))$ $=\mathbb{E}_{q_{i}}(\nabla_{\bm{\theta}}\log p(\mathbf{y}_{i}\mid\bm{\theta})),$ (S5)
$\nabla_{\mathbf{\Sigma}_{i}}\mathbb{E}_{q_{i}}(\log p(\mathbf{y}_{i}\mid\bm{\theta}))$ $=\frac{1}{2}\mathbb{E}_{q_{i}}(\nabla_{\bm{\theta}}^{2}\log p(\mathbf{y}_{i}\mid\bm{\theta})).$ (S6)
Thus the updates (S3) and (S4) become
$\bm{\mu}_{i}$ $=\bm{\mu}_{i-1}+\mathbf{\Sigma}_{i-1}\mathbb{E}_{q_{i}}(\nabla_{\bm{\theta}}\log p(\mathbf{y}_{i}\mid\bm{\theta})),$ (S7)
$\mathbf{\Sigma}_{i}^{-1}$ $=\mathbf{\Sigma}_{i-1}^{-1}-\mathbb{E}_{q_{i}}(\nabla_{\bm{\theta}}^{2}\log p(\mathbf{y}_{i}\mid\bm{\theta})).$ (S8)
These updates are implicit as they require the evaluation of expectations with respect to $q_{i}(\bm{\theta})$. Under the assumption that $q_{i}(\bm{\theta})$ is close to $q_{i-1}(\bm{\theta})$, Lambert et al., (2022) propose replacing $q_{i}(\bm{\theta})$ with $q_{i-1}(\bm{\theta})$ in (S7) and (S8), and replacing $\mathbf{\Sigma}_{i-1}$ with $\mathbf{\Sigma}_{i}$ on the right hand side of (S7), to yield an explicit scheme
$\bm{\mu}_{i}$ $=\bm{\mu}_{i-1}+\mathbf{\Sigma}_{i}\mathbb{E}_{q_{i-1}}(\nabla_{\bm{\theta}}\log p(\mathbf{y}_{i}\mid\bm{\theta})),$ (S9)
$\mathbf{\Sigma}_{i}^{-1}$ $=\mathbf{\Sigma}_{i-1}^{-1}-\mathbb{E}_{q_{i-1}}(\nabla_{\bm{\theta}}^{2}\log p(\mathbf{y}_{i}\mid\bm{\theta})).$ (S10)
Equations (S9) and (S10) form the so-called R-VGA algorithm of Lambert et al., (2022).
[…]
S2 Additional details on the SV model
In this section, we follow Ruiz, (1994) and show that the terms $\{\xi_{t}\}_{t=1}^{T}$ in the log-squared transformation (27) in the main paper have mean zero and variance $\pi^{2}/2$.
Since $\epsilon_{t}\sim N(0,1)$, we have $\epsilon_{t}^{2}\sim\chi^{2}_{1}$, that is, $\epsilon^{2}_{t}$ is gamma distributed with shape and scale parameters equal to $1/2$ and $2$, respectively. Therefore, $\log(\epsilon_{t}^{2})$ has mean $\psi(1/2)+\log(2)$, where $\psi(\cdot)$ denotes the digamma function (the first derivative of the logarithm of the gamma function), and variance $\psi^{(1)}\left(\frac{1}{2}\right)=\frac{\pi^{2}}{2}$, where $\psi^{(1)}(\cdot)$ is the trigamma function (the second derivative of the logarithm of the gamma function; see, e.g., Abramowitz and Stegun, (1968)). Thus
$\mathbb{E}(\xi_{t})=\mathbb{E}\left(\log(\epsilon_{t}^{2})-\mathbb{E}(\log(\epsilon_{t}^{2}))\right)=0,$
and
$\operatorname{Var}(\xi_{t})=\operatorname{Var}\left(\log(\epsilon_{t}^{2})-\mathbb{E}(\log(\epsilon_{t}^{2}))\right)=\operatorname{Var}\left(\log(\epsilon_{t}^{2})\right)=\frac{\pi^{2}}{2}.$
S3 Experiments involving different blocking strategies
In this section, we discuss the tuning parameters in Algorithm 3: the choice of block length $B$ and the number of frequencies to be processed individually before blocking, $\tilde{n}$. We base our discussion on inspection of the periodogram.
Recall that the periodogram is an estimate of the spectral density and can be viewed as a measure of the relative contributions of the frequencies to the total variance of the process (Percival and Walden,, 1993). For example, Figure S1 shows the periodograms for the linear Gaussian SSM in Section 3.1 and the univariate SV model in Section 3.2. This figure shows that in general, power decreases as frequency increases. Frequencies that carry less power tend to have less effect on the trajectory of the variational mean; see, for example, the trajectories of the parameters of the linear Gaussian model in Figure S2. The trajectory of $\phi$ moves towards the true parameter value within the first few frequencies, with only minor adjustments afterwards. The trajectories for $\sigma_{\eta}$ and $\sigma_{\epsilon}$ behave in a similar manner, but with more gradual changes. This behaviour can also be observed in Figure S3, in which the trajectories of both parameters in the univariate SV model exhibit large changes in the first few iterations, then stabilise afterwards. Higher frequencies that have little effect on the variational mean can thus be processed together in “blocks”. The R-VGA-Whittle updates can then be made based on the Whittle log-likelihood for a “block” of frequencies, which we compute as the sum of the individual log-likelihood contribution evaluated at each frequency within that block.
(a) Linear Gaussian model.
(b) Univariate SV model.
Figure S1: Plot of raw periodograms of the data (black) and smoothed periodograms (light red) for two models. Vertical lines show the frequency cutoffs at half-power (3dB, red dashed line), one-fifth-power (7dB, blue dotted line) and one-tenth-power (10dB, purple dash-dotted line).
Figure S2: Trajectories of the variational mean for the parameters in the LGSS model estimated using damped R-VGA-Whittle (Algorithm 2) with no blocking. The true parameter values are marked with dashed lines.
Figure S3: Trajectories of the variational mean for the parameters in the univariate SV model estimated using damped R-VGA-Whittle (Algorithm 2) with no blocking. The true parameter values are marked with dashed lines.
It is important to choose an appropriate point to start blocking, as well as an appropriate length for each block of frequencies. To determine the frequency at which to begin blocking, we borrow a concept from physics and electrical engineering called the 3dB frequency cutoff, which is the frequency on the power spectrum at which the power drops to approximately half the maximum power of the spectrum. The value 3 comes from the fact that, on the decibel scale, a drop from a power level $P_{1}$ to another level $P_{2}=\frac{1}{2}P_{1}$ results in a drop of 3 decibels. The change in decibels is calculated as
$\Delta\textrm{dB}=10\log_{10}\frac{P_{2}}{P_{1}}=10\log_{10}\left(\frac{1}{2}\right)=-3.0103.$
We employ a similar idea here, where we define the “cutoff” frequency as the frequency at which the power drops to half of the maximum power in a smoothed version of the periodogram. Specifically, we seek the frequency $\omega^{*}$ such that $\tilde{I}(\omega^{*})=1/2\max(\tilde{I}(\omega))$, where $\tilde{I}(\omega)$ is the periodogram smoothed using Welch’s method (Welch,, 1967). Briefly, Welch’s method works by dividing the frequencies into equal segments, computing the periodogram for each segment, and then averaging over these segments to obtain a smoother periodogram with lower variance. We also experiment with other cutoffs, such as the frequencies at which the power are one-fifth and one-tenth of the maximum power in the smoothed periodogram, which correspond to a drop of approximately 7dB and 10dB, respectively. We mark these cutoff frequencies in Figure S1.
Using each of the 3dB, 7dB and 10dB cutoffs, we run R-VGA-Whittle with no blocking, and then with varying block sizes ($B=10,50,100,300,500$, and $1000$ frequencies), and examine the posterior densities of the parameters under these different settings. For all runs, we keep the settings for other tuning parameters the same as in Section 3 and use $S=1000$ Monte Carlo samples, $n_{damp}=5$, and $D=100$ damping steps. For the linear Gaussian SSM, we plot the R-VGA-Whittle posteriors for these different cutoffs and block sizes in Figures 4(a), 4(b) and 4(c). The corresponding plots for the univariate SV model are shown in Figures 5(a), 5(b), and 5(c). The plots for the bivariate SV model are similar to those from the univariate SV model and not shown.
We find that for the linear Gaussian SSM, block sizes of $100$ and below result in R-VGA-Whittle posterior densities that are very similar to each other, similar to the posterior densities obtained without blocking, and also similar to the posterior densities from HMC-Whittle and HMC-exact. For block sizes of $300$ and $500$, the posterior densities of the parameters $\sigma_{\eta}$ and $\sigma_{\epsilon}$ begin to differ slightly compared to the posterior densities obtained without blocking. When the block size is $1000$, the posterior densities for all three parameters become noticeably biased, although this bias reduces when the frequency cutoff point increases, as expected. However, for the SV model, the R-VGA-Whittle posterior densities are highly similar for all block sizes. To achieve the best reduction in computing time while retaining estimation accuracy, we choose to begin blocking at the earliest cutoff point we consider in this experiment (3dB), with the block size set to 100.
(a) Blocking begins at the 3dB cutoff point.
(b) Blocking begins at the 7dB cutoff point.
(c) Blocking begins at the 10dB cutoff point.
Figure S4: Posterior densities of the parameters of the linear Gaussian SSM estimated using R-VGA-Whittle without blocking (denoted as “None”) and with blocking starting at different cutoff points for several block sizes (coloured lines), HMC-Whittle (dashed lines), and HMC-exact (dotted lines). True parameter values are denoted using vertical dashed lines.
(a) Blocking begins at the 3dB cutoff point.
(b) Blocking begins at the 7dB cutoff point.
(c) Blocking begins at the 10dB cutoff point.
Figure S5: Posterior densities of the parameters of the univariate SV model estimated using R-VGA-Whittle without blocking (denoted as “None”) and with blocking starting at different cutoff points for several block sizes (coloured lines), HMC-Whittle (dashed lines), and HMC-exact (dotted lines). True parameter values are denoted using vertical dashed lines.
S4 Variance of the R-VGA-Whittle posterior densities for various Monte Carlo sample sizes
In this section, we study the robustness of R-VGA-Whittle posterior distributions to different values of $S$, the Monte Carlo sample size used to approximate the expectations of the gradient and Hessian in Algorithm 2. We consider sample sizes $S=100,S=500$ and $S=1000$. We apply blocking as in Algorithm 3, and fix the block length to $B=100$. We set $\tilde{n}=176$ for the linear Gaussian SSM, and $\tilde{n}=75$ for the univariate SV model, which corresponds to the 3dB cutoff point as discussed in Section S3.
Figures 6(a)6(c) show the posterior densities of the parameters of the linear Gaussian SSM from 10 runs of R-VGA-Whittle with $S=100$, $S=500$, and $S=1000$, respectively, while Figures 7(a)7(c) show the corresponding plots for the SV model. As the Monte Carlo sample size increases, the variance in the posterior densities across different runs decreases. When $S=1000$, the variance across different runs is significantly reduced, so we choose this as our Monte Carlo sample size for all examples in Section 3 of the main paper.
(a) $S=100$.
(b) $S=500$.
(c) $S=1000$.
Figure S6: Posterior densities of the parameters of the linear Gaussian SSM from 10 runs of R-VGA-Whittle (coloured solid lines) with different Monte Carlo sample sizes $S$, from a single run of HMC-Whittle (dashed lines), and from a single run of HMC-exact (dotted lines). True parameter values are marked with dashed lines.
(a) $S=100$.
(b) $S=500$.
(c) $S=1000$.
Figure S7: Posterior densities of the parameters of the univariate SV model from 10 runs of R-VGA-Whittle (coloured solid lines) with different Monte Carlo sample sizes $S$, from a single run of HMC-Whittle (dashed lines), and from a single run of HMC-exact (dotted lines). True parameter values are marked with dashed lines.
S5 Trace plots
This section shows the HMC-exact and HMC-Whittle trace plots and effective sample size (ESS) for the parameters in all the models we consider. Figure S8 shows trace plots for the linear Gaussian SSM, Figures S9 and S10 show trace plots for the univariate and bivariate SV models with simulated data, and Figures S11 and S12 show trace plots for the univariate and bivariate SV models with real data.
In general, we find that HMC-Whittle tends to produce posterior samples with much lower correlation than HMC-exact. For the same number of posterior samples, HMC-Whittle has higher ESS than HMC-exact, and the ESS of HMC-Whittle is high for all paramaters, while HMC-exact tends to produce higher ESS for the autoregressive parameters ($\phi$ for univariate examples and $\mathbf{\Phi}$ for bivariate examples) than for the variance parameters ($\sigma_{\eta}$ and $\sigma_{\epsilon}$ for the linear Gaussian model, $\sigma_{\eta}$ for the univariate SV model, and $\mathbf{\Sigma}_{\eta}$ for the bivariate SV model). This is because HMC-Whittle only targets the posterior density of the parameters, whereas HMC-exact targets the joint posterior density of the parameters and states, which are high dimensional. See Table S1 for the ESS corresponding to each parameter in our examples.
Example Parameter ESS (HMC-exact) ESS (HMC-Whittle)
Linear Gaussian $\phi$ 2452.8033 8884.775
Linear Gaussian $\sigma_{\eta}$ 846.3039 7933.656
Linear Gaussian $\sigma_{\epsilon}$ 697.3547 8163.235
Univariate SV (sim. data) $\phi$ 230.00751 6854.622
Univariate SV (sim. data) $\sigma_{\eta}$ 87.94789 6949.066
Bivariate SV (sim. data) $\Phi_{11}$ 814.3441 13064.54
Bivariate SV (sim. data) $\Phi_{22}$ 282.3187 12967.59
Bivariate SV (sim. data) $\Sigma_{\eta_{11}}$ 346.6532 13382.68
Bivariate SV (sim. data) $\Sigma_{\eta_{21}}$ 440.2904 16141.23
Bivariate SV (sim. data) $\Sigma_{\eta_{22}}$ 138.6995 12034.28
Univariate SV (real data) $\phi$ 304.6549 5920.983
Univariate SV (real data) $\sigma_{\eta}$ 118.8147 5980.517
Bivariate SV (real data) $\Phi_{11}$ 730.7380 9058.774
Bivariate SV (real data) $\Phi_{22}$ 796.0991 9242.738
Bivariate SV (real data) $\Sigma_{\eta_{11}}$ 206.4234 11170.016
Bivariate SV (real data) $\Sigma_{\eta_{21}}$ 161.9922 9989.325
Bivariate SV (real data) $\Sigma_{\eta_{22}}$ 245.5852 11127.356
Table S1: The ESS obtained for all parameters considered in the examples described in the main text. The ESS is shown for both HMC-exact and HMC-Whittle from 20000 posterior samples.
(a) HMC-exact trace plots (samples not thinned).
(b) HMC-Whittle trace plots (samples not thinned).
Figure S8: Trace plots for the linear Gaussian SSM with simulated data. Red dashed lines denote the true parameter values.
(a) HMC-exact trace plots (samples not thinned).
(b) HMC-exact trace plots (samples thinned by a factor of 100).
(c) HMC-Whittle trace plots (samples not thinned).
Figure S9: Trace plots for the univariate SV model with simulated data. Red dashed lines denote the true parameter values.
(a) HMC-exact trace plots (samples not thinned).
(b) HMC-exact trace plots (samples thinned by a factor of 50).
(c) HMC-Whittle trace plots (samples not thinned).
Figure S10: Trace plots for the bivariate SV model with simulated data. Red dashed lines denote the true parameter values.
(a) HMC-exact trace plots (samples not thinned).
(b) HMC-exact trace plots (samples thinned by a factor of 100).
(c) HMC-Whittle trace plots (samples not thinned).
Figure S11: Trace plots for the univariate SV model applied to exchange rate data.
(a) HMC-exact trace plots (samples not thinned).
(b) HMC-exact trace plots (samples thinned by a factor of 50).
(c) HMC-Whittle trace plots (samples not thinned).
Figure S12: Trace plots for the bivariate SV model applied to exchange rate data.